OpenStudy (anonymous):

Exponential Equations Solve 12^x-2=20 Round to the nearest ten-thousandth x=__________

OpenStudy (anonymous):

is it \[\large 12^{x-2}=20\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

use the change of base formula \[\large b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\]

OpenStudy (anonymous):

so what would the answer be?

OpenStudy (anonymous):

whatever you get when you replace \(A\) by \(20\), \(b\) by \(12\) and \(x\)by \(x-2\) in the above change of base formula you will need a calculator to compute the number as a decimal

OpenStudy (anonymous):

ok i will plug it in, thank you

OpenStudy (anonymous):

do you happen to have a calculator, im using a pelletty free one on the computer.